wiki:UC20210601

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 06/01/2021


User feedback and comments from institutional representatives

  • NAPG - no report
  • Washington - nothing to report
  • Georgia State - nothing to report
  • FGCU - nothing to report
  • Seattle - nothing to report
  • Oklahoma - nothing to report (status provided by email)
  • Colorado - question from user: the Q3 proposal call said KOSMOS might be available, requesting info re: timeline (see below for KOSMOS status update)
  • Wyoming - nothing to report
  • NMSU - nothing to report (status provided by email)
  • JHU - nothing to report
  • BYU - nothing to report
  • UVa - nothing to report

Telescope and Instruments Report

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 5/01/2021 - 5/28/2021

0) Overview

APO transitioned to a 75% staffing limit at the beginning of May, when Otero County moved from the Green tier to the Turquoise tier in the New Mexico Department of Health COVID framework. The Turquoise level allows APO to have essentially full staff.

The first visitors on site at APO in over a year took place in May as well; a team from UVa was on site working on the APOGEE instrument.

Weather for the month of May has been warming, with occasional monsoon weather.

The new 3.5m Observer started on 5/24. Amanda Townsend will be working in tandem with the other observing specialists while she is being trained.

There has been a consistent communications issue affecting remote observers connecting to APO this month. There are dramatic slowdowns of the connection to APO until about midnight/1am MDT. This was initially referred to as the “Comcast” problem because everyone affected by it had Comcast/Xfinity internet service. However, we are now seeing that it affects more than that group. The issue has been traced to L3 communications (a major ISP) having an problem with slow connections between some ISP’s. APO’s ISP has a ticket open with L3 and is working the issue. There was a small change put in place by our ISP that improved things from an “unusable” state to “annoying but workable,” as reported by Dr. Derek Buzasi.

1) Telescope

The telescope is operational, however, it is experiencing seasonal harmonic oscillations and audio ringing for especially the altitude axis. Some motion control parameters have been adjusted to minimize the issue, however it hasn’t gone away completely. We are monitoring the situation and will adjust parameters as needed.

2) Instruments

DIS: Scattered light is still quite good for both red and blue. The red camera temperatures are elevated, which may necessitate that darks be taken during cals. The observing staff at APO is going to confirm that the darks are flat (consistent structure if any) and report to the users if the dark current can just be easily subtracted out.

A failure of the detent system has made stability of CWL calibrations difficult and the current will move with instrument rotations. The instrument may be unusable except at 0 degrees rotator position. The engineering staff is troubleshooting this.

TripleSpec and Agile: powered on, cold and operational.

Echelle: The ion pump and vacuum servicing was performed this past month. However, within 48 hours of returning the instrument to service the interorder light ratios dropped to the point where the instrument was unusable for science. The CCD reservoir was warmed, while keeping the shield cold, for 4 days. This was partially successful. Within another 48 hours the IOL ratios began slowly dropping. We are monitoring the situation daily. We have seen this behavior in the past before a leveling off as well as a continual drop.

NICFPS: The instrument is performing nominally with the occasional need to reset the controller due to corrupted images. We recommend users watch their images carefully for scrambled, excessive noise or blank frames; if any are seen they should inform the Observing Specialist. Recovery back to operations is generally quick.

ARCTIC: The instrument is performing nominally. Ion pump currents are elevated and are indicating that a replacement is needed soon. We are scheduling vacuum servicing the week before summer shutdown and it should be available about the same time as return to science.

3) ARCSAT There was a recent attempt to collect more data with the Shelyak. However, network problems, weather, and then a failed cooler on the science camera prevented any data from being collected. Temperature cooling has been repaired in house. Bill is working with Diffraction Limited on the bugs and features in MaxIm DL. Simultaneous observations were obtained during 5/27 engineering with Shelyak and Echelle, a milestone for Shelyak commissioning.


Additional telescope and instrument discussion

Over the last month the observatory has returned to the 75% staffing level; people are coming back to work but maintaining social distance. First visitors to the site in over a year arrived in May. The weather is warming, which means the moths are back. The new Observing Specialist, Amanda Townsend, started last week. There is an ongoing communications issue between the observatory and outside world (see above). Derek Buzasi's recent experience was that remote observing worked OK for the 3.5m but less well for ARCSAT, likely because there is lots more info going back and forth. UC reps: warn your users of potentially slow connection speeds to the observatory. The obs-specs can help with target acquisitions etc if users' connections are slow. The telescope is experiencing a seasonal harmonic issue that causes the altitude drive to oscillate. The engineering staff made some tweaks so the situation should be improving. The obs-specs have a procedure to move the telescope to cancel the oscillations if they appear. This boils down to a choice between making things good at tracking speed or at slew speed; things are better at tracking and the error shows up during slew. Users should be alert for the possibility of pointing errors. The DIS temperature dropped, which is good, but currently there is an issue with the detent system, affecting the turret motor that positions the turret for the desired grating. The engineering staff is working on it soon. ARCES had its vacuum serving in May and it struggled to come back to optimum performance afterwards. Bill worked on it. The inter-order light is worse than last year but not unusable; it might be a problem for some kinds of science (bright emission lines could bleed into adjacent orders, or problems identifying as many ThAr lines in blue as in previous years). UC reps: warn users to be aware of this and inform us if they are having trouble achieving their science goals. Finally, for NICFPS, UC reps should warn NICFPS users to look at their images and let the obs-spec know if they are getting a blank frame or jumbled signal. For ARCSAT, we recently achieved joint observations with the 3.5m and Shelyak, which will be important for assessing the Shelyak's performance. There was a cooler issue that was fixed. Network and software issues continue to be a problem; hopefully there will be a new software update soon.


Discussion of summer shutdown status

The primary mirror will not be realuminized this summer as Kitt Peak couldn't work us into their schedule until October, which is not viable for us. Thus, we will have a shorter shutdown this summer of approximately 2 weeks, from 19-29 July. The return to science will be contingent upon sufficiently good weather toward the end of shutdown to conduct on-sky engineering, so the shutdown end date is subject to change. ARCTIC will be removed from service the week before shutdown begins for routine maintenance.


KOSMOS Update

Although KOSMOS may be available in Q3 users should consider it in shared risk mode at that time. The plan is to bring it online in the TEG lab at UW within the next 1-2 weeks. COVID restrictions at UW have eased and they are nearly back to full staffing in the lab. They are coordinating with the site staff regarding shipping the instrument to APO.


ARCTIC Diffuser Rotator Update

The part was sent back to UW for some additional engineering. They are making some small changes before sending it back.


Q2 3.5m scheduling

There is a fair number of open slots remaining in June. Users are reminded check the Q2 schedule (both DD01 and OPEN time) and to follow the standard procedure of emailing requests for this time to Ben, Russet, Nancy, and their institutional scheduler.


Q2 ARCSAT scheduling

There is no open time available in Q2. The call for Q3 proposals will be coming out soon. [Note added in proof: the call for ARCSAT proposals was sent on 06/11/2021.] There is a week at the end of June that straddles Q2 and Q3 so users who wish to request that time should submit their proposals ASAP, even before the nominal deadline of June 24.


Q3 3.5m scheduling

Russet has started on the schedule; it's too soon to say whether there are any major conflicts or oversubscriptions. It looks like there is a request for one class and one non-student training session. [Note added in proof: the Q3 schedule is now published and available online.]


Quarterly Password Notifications

Now that people are returning to work in their offices more, if those who requested that the quarterly passwords get sent to their homes would now like them to be sent to their offices, please let Nancy know.


3.5m Science Symposium

The virtual APO science symposium will be held on 26-28 July. This symposium is intended to showcase results from observations made with the 3.5m and ARCSAT telescopes. Students are particularly encouraged to present their research. A website and registration page are in the works and will be announced within the next couple of weeks.


UC Meeting Schedule for July-December

Our Users' Committee meetings are scheduled through June. Nancy proposed leaving it at the same day and time (first Tuesday of the month at 10:30 am Mountain Time). If anyone has any conflicts please let Nancy know.


ACTION ITEMS

Open action items from previous meetings:

  • UC reps - remind users of deadline for requesting non-student training session in the upcoming quarter. Status: CLOSED.

New action items from this meeting:

  • UC reps - warn your users of the following potential issues: (1) slow connection speeds to the observatory, (2) pointing errors after a slew due to oscillations in the altitude drive, (3) increased inter-order light on ARCES, (4) check NICFPS images and let the obs-spec know if they are getting a blank frame or jumbled signal so they can reset some of the electronics.

All Other Business

None.


Next meeting

The next meeting will be on Tuesday July 6 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.